OpenStudy (anonymous):

Which answer correctly estimates the product of 20.21 × 999.801? A. about 20,000 B. about 2,000 C. about 200 D. about 20

OpenStudy (anonymous):

@Utterly_Confuzzled

OpenStudy (anonymous):

help plz

OpenStudy (utterly_confuzzled):

round 20.21 to 20 and 999.801 to 1000. 20*1000=??

OpenStudy (lizz123):

Round \[20.01\rightarrow 20.00\] round \[999.801\rightarrow1000\]

OpenStudy (lizz123):

Now you have 20.00*1000=?

OpenStudy (anonymous):

a

OpenStudy (lizz123):

Solve the problem and you will get your answer

OpenStudy (anonymous):

i have another one wait plz

OpenStudy (lizz123):

correct I got that too

OpenStudy (utterly_confuzzled):

Correct, A is the right answer

OpenStudy (anonymous):

Which answer correctly estimates the quotient of 102.02 ÷ 4.011? A. about 25 B. about 250 C. about 2,500 D. about 25,000

OpenStudy (lizz123):

\[102.02\rightarrow100.00\] \[4.011\rightarrow4.000\]

OpenStudy (lizz123):

Divide that

OpenStudy (anonymous):

25

OpenStudy (lizz123):

Yes, you are correct. Do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (lizz123):

what is it

OpenStudy (anonymous):

Estimate the product of 15.217 × 9.54. Round each number to the nearest whole number. about ____________

OpenStudy (anonymous):

no

OpenStudy (utterly_confuzzled):

15.217 rounds to 15 9.54 rounds to 10 15x10 = ???

OpenStudy (anonymous):

150

OpenStudy (anonymous):

Estimate the quotient of 719.711 ÷ 8.75. Round the dividend and divisor to the nearest whole number. about

OpenStudy (anonymous):

help me plz thank u

OpenStudy (anonymous):

last one thats it

OpenStudy (utterly_confuzzled):

719.711 rounds to 720 8.75 rounds to 9 720/9 =????

OpenStudy (anonymous):

80

OpenStudy (anonymous):

thank u bye